# Existence results for a biofilm free-boundary problem with dominant detachment

### v1

@@ Abstract
This work addresses the existene and uniqueness of a Wanner-Gujer free boundary problem modeling biofilms with a prevailing detachment. This result extends the some works from tumor gowth modeling as well as those from the biofilm modeling field. Besides the local existence and uniqueness, the continuous dependence on initial and boundary data and the global existence is also given. The local existence is obtained by using fixed point arguments combined with semigroup theory while the global existence is deduced using invariance regions and energy estimates.
@@
Biofilms , Free boundary problem , Semigroup , Anisotropic spaces , Hyperbolic and Parabolic PDE , Energy estimates MSC 35A20 , 35A08 , 35A07 , 35A22 , 35C15 , 35R35
journal: Applied Mathematics letters
[…]
organization=Départment de Mathématique-Physique, Institut Supérieur Pédagogique de Bukavu,addressline=30 Avenue Kibombo, city=Bukavu, postcode=854, country=Democratic Republic of Congo
1 Introduction
@@ 2 Reformualtion of the problem
Define the following vector-valued functions : $\boldsymbol{C}=(C_{1},\cdots,C_{m})^{T},\boldsymbol{Y}=(Y_{1},\cdots,Y_{n})^{T},$ $\boldsymbol{F}=(F_{1},\cdots,F_{n})^{T},\boldsymbol{H}=(H_{1},\cdots,H_{m})^{T},\boldsymbol{B}(t)=B(t)\boldsymbol{I}_{m}$ where $B(t)=zv(1,t)$ and $\boldsymbol{I}_{m}$ the identity matrix of order $m.$ This system is equivalent to equations (1)–(3) and will be the focus of the rest of the work. We will first give primary lemmas which will be used in the proof of the main results.
3 Primary lemmas and local existence
@@
Proof. The proof is divided into several steps, the first of which is the parabolic lifiting of boundary data. In fact, since $\boldsymbol{\Psi}\in W^{1-\frac{1}{2p},p}(0,T_{1})^{m}$, by the parabolic trace theorem (see for example [26, 27]), there exists $\tilde{\boldsymbol{\Psi}}\in W^{2,1}_{p}(\Omega\times(0,T_{1}))^{m}$ such that:
  • 1.
@@
with $\mathcal{F}(t,\phi)(z)=\boldsymbol{H}(\boldsymbol{Y}(z,t),\phi(z)+\tilde{\boldsymbol{\Psi}}(z,t))+(\boldsymbol{B}(t)\partial_{z}+\partial_{zz}-\partial_{t})\tilde{\boldsymbol{\Psi}}(z,t)$. Let $\mathcal{A}_{0}\phi=\partial_{zz}\phi$ with domain $D$. On one hand, by [28, Theorem 2.7, Chap.7] and [29, Chap. 8-9], $\mathcal{A}_{0}$ is strongly elliptic and generates an analytic semigroup on $\mathbb{X}$.
@@
The estimate follows from the fixed point property and the linear estimates. The second-order compatibility condition ensures that the solution maintains the $W^{2,1}_{p}$ regularity up to $t=0$. $\square$
Lemma 3.2
@@
Proof. Since $v(1,t)$ depends only on $t$, we define the characteristic curve
@@
so the solution remains nonnegative. This completes the proof. $\square$
For the local existence we will need the following assumptions in addition to those admititted in Lemmas 3.1 and 3.2.
@@
Moreover there exist constants $K_{1},K_{2}>0$ depending only on the data norms $\|\varphi\|_{C^{1}},\|\theta\|_{W^{2-2/p,p}},\|\psi\|_{W^{1,p}},R_{0}$, the Lipschitz constants, and $p$, such that for $0\leq t\leq T^{\ast}$:
@@
Proof. The proof relies again on a fixed point argument on a small time interval. We split the construction into maps solved by the two given lemmas, $v$ and for $R$. By fixing $T>0$ to be chosen and we define the Banach space
@@
Finally, standard continuation arguments apply: the solution can be extended past $T^{\ast}$ as long as the norms appearing in $\|\boldsymbol{C}\|_{W^{2,1}_{p}}+\|\boldsymbol{Y}\|_{C^{1}}+\|R\|_{C^{1}}$ remain finite. This finishes the proof. $\square$
4 Global existence via invariance regions and energy estimates
@@
This inequality (27) is a dissipation condition which ensures that reaction terms do not allow arbitrary growth and provide a restoring effect for large norms. To prove this system we use the local existence result and the continuation criterion then derive a priori energy estimates which are uniform on finite time intervals and in fact globally bounded. We conclude the proof by using these estimates to prevent the norms that control continuation from blowing up; hence extend the solution globally.
Proof. By the local well-posedness result in Theorem 3.3) we only need to obtain a priori bounds preventing blow-up of the estimates obtained earlier. In addition, since each $f_{i}$ and $h_{j}$ are quasi-positive (hypothesis 3)) and the transport and diffusion operators preserve sign under the given boundary conditions, standard comparison-principle arguments yield
@@
Since all continuation norms remain bounded for all finite times, the local solution can be extended step by step to a global solution on $0,\infty)$. The uniform-in-time a priori bounds (27) prevent finite-time blow-up and guarantee classical regularity persists for all times. $\square$

### v2

@@ Abstract
This work addresses the existence and uniqueness of a Wanner-Gujer free-
boundary problem that models biofilms under conditions of prevailing detachment.
This result significantly extends previous findings in both tumor growth modeling and
the biofilm modeling field.
Besides establishing local existence and uniqueness, we also prove the continuous de-
pendence of the solution on initial and boundary data. Furthermore, global existence
is deduced using a combination of invariance regions and energy estimates. The proof
for local existence is obtained by utilizing fixed point arguments combined with
semigroup theory.
@@
Biofilms, Free boundary problems, Semigroups, Anisotropic spaces,
Hyperbolic and Parabolic PDE, Energy estimates
1991 Mathematics Subject Classification:
35A20 ;35A08 ; 35A07 ;35A22 ;35C15 ;35R35
1. Introduction
@@
In the following pages the work is organised as follows: the second section introduces the notations used and presents a reformuation of the free boundary problem into an initial value problem. In the third section we present primary lemmans and the main result for local existence. Finally, the last section is about the global existence by using the region invariance annd energy estimates.
2. Notations and reformualtion of the problem
@@
Define the following vector-valued functions : $\boldsymbol{C}=(C_{1},\cdots,C_{m})^{T},\boldsymbol{Y}=(Y_{1},\cdots,Y_{n})^{T},$ $\boldsymbol{F}=(F_{1},\cdots,F_{n})^{T},$
$\boldsymbol{H}=(H_{1},\cdots,H_{m})^{T},\boldsymbol{B}(t)=B(t)\boldsymbol{I}_{m}$ where $B(t)=zv(1,t)$ and $\boldsymbol{I}_{m}$ the identity matrix of order $m.$ This system is equivalent to equations (1)–(3) and will be the focus of the rest of the work. We will first give primary lemmas which will be used in the proof of the main results.
3. Primary lemmas and local existence
@@
Phys. Rev. E The proof is divided into several steps, the first of which is the parabolic lifiting of boundary data. In fact, since $\boldsymbol{\Psi}\in W^{1-\frac{1}{2p},p}(0,T_{1})^{m}$, by the parabolic trace theorem (see for example [2, 4]), there exists $\tilde{\boldsymbol{\Psi}}\in W^{2,1}_{p}(\Omega\times(0,T_{1}))^{m}$ such that:
@@
with $\mathcal{F}(t,\phi)(z)=\boldsymbol{H}(\boldsymbol{Y}(z,t),\phi(z)+\tilde{\boldsymbol{\Psi}}(z,t))+(\boldsymbol{B}(t)\partial_{z}+\partial_{zz}-\partial_{t})\tilde{\boldsymbol{\Psi}}(z,t)$. Let $\mathcal{A}_{0}\phi=\partial_{zz}\phi$ with domain $D$. On one hand, by [17, Theorem 2.7, Chap.7] and [20, Chap. 8-9], $\mathcal{A}_{0}$ is strongly elliptic and generates an analytic semigroup on $\mathbb{X}$.
@@
The estimate follows from the fixed point property and the linear estimates. The second-order compatibility condition ensures that the solution maintains the $W^{2,1}_{p}$ regularity up to $t=0$.
Lemma 3.2.
@@
Phys. Rev. E Since $v(1,t)$ depends only on $t$, we define the characteristic curve
@@
so the solution remains nonnegative. This completes the proof. For the local existence we will need the following assumptions in addition to those admititted in Lemmas 3.1 and 3.2.
@@
Moreover there exist constants $K_{1},K_{2}>0$ depending only on the data norms $\|\varphi\|_{C^{1}},$
$\|\theta\|_{W^{2-2/p,p}},\|\psi\|_{W^{1,p}},R_{0}$, the Lipschitz constants, and $p$, such that for $0\leq t\leq T^{\ast}$:
@@
Phys. Rev. E The proof relies again on a fixed point argument on a small time interval. We split the construction into maps solved by the two given lemmas, $v$ and for $R$. By fixing $T>0$ to be chosen and we define the Banach space
@@
Finally, standard continuation arguments apply: the solution can be extended past $T^{\ast}$ as long as the norms appearing in $\|\boldsymbol{C}\|_{W^{2,1}_{p}}+\|\boldsymbol{Y}\|_{C^{1}}+\|R\|_{C^{1}}$ remain finite. This finishes the proof.
4. Global existence via invariance regions and energy estimates
@@
This inequality (27) is a dissipation condition which ensures that reaction terms do not allow arbitrary growth and provide a restoring effect for large norms. To prove this system we use the local existence result and the continuation criterion then derive a priori energy estimates which are uniform on finite time intervals and in fact globally bounded. We conclude the proof by using these estimates to prevent the norms that control continuation from blowing up; hence extend the solution globally. Phys. Rev. E By the local well-posedness result in Theorem 3.3) we only need to obtain a priori bounds preventing blow-up of the estimates obtained earlier. In addition, since each $f_{i}$ and $h_{j}$ are quasi-positive (hypothesis 3)) and the transport and diffusion operators preserve sign under the given boundary conditions, standard comparison-principle arguments yield
@@
Since all continuation norms remain bounded for all finite times, the local solution can be extended step by step to a global solution on $0,\infty)$. The uniform-in-time a priori bounds (27) prevent finite-time blow-up and guarantee classical regularity persists for all times.
5. Statements and Declarations
Competing Interests: The author declares no competing interests.